# Practical Collaborative Perception: A Framework for Asynchronous and Multi-Agent 3D Object Detection

### v1

@@ III-A Improving Single-Vehicle Perception
As mentioned in Sec.I, the accuracy of single-vehicle detection can be significantly improved by using the concatenation point cloud sequences as inputs. However, such concatenation is done by the Ego Motion Compensation method which only accounts for the motion of the ego vehicle resulting in the shadow effect (Fig.1) in the concatenated point cloud. This effect, which is a misalignment in the 3D space between objects’ points and their locations, leads to a misalignment in feature spaces (e.g., BEV representations), thus degrading detection accuracy for dynamic objects. In this section, we recall the approach we use in [1] to handle this effect and present the extensions we make to it.
III-A1 Shadow Effect: Cause and Rectification
@@
Let $\mathcal{P}^{t}=\left\{\mathbf{p}_{j}^{t}=[x,y,z,r,\Delta]~{}|~{}j=1,\dots,N\right\}$ denote a point cloud collected by the ego vehicle at time step $t$. Here, each point $\mathbf{p}_{j}$ has two features namely reflectant $r$ and time-lag $\Delta$ with respect to a predefined time step. The point cloud is expressed in the ego vehicle frame $\mathcal{E}(t)$ measured with respect to a global frame $\mathcal{G}$. To concatenate or a point cloud sequence $\mathcal{S}=\left\{\mathcal{P}^{t-K},\mathcal{P}^{t-K+1},\dots,\mathcal{P}^{t}\right\}$ of length $K+1$, each point of point cloud $\mathbf{P}^{t-\Delta t}$ is transformed to the global frame using the ego vehicle pose at the same time step as in (1), hence the method’s name Ego Motion Compensation (EMC).
@@
Here, ${}^{d}\mathbf{T}_{s}\in\mathbf{SE(3)}$ represents the rigid transformation that maps points in frame $s$ to frame $d$. EMC undoes the motion of the ego vehicle; however, it fails to account for objects’ motion. As a result, the appearance of a dynamic object in the concatenated point cloud comprises several instances, each corresponding to the object’s poses in the global frame at a particular time step, as in Fig.1.
@@ III-A2 Learning to Rectify the Shadow Effect
In [1], we propose a drop-in module, named Aligner, for 3D object detection models that create a Bird-Eye View (BEV) representation, $\mathbf{B}\in\mathbb{R}^{C\times H\times W}$, of the scene as an intermediate output. Our module first computes the feature of $\mathbf{f}_{i}\in\mathbb{R}^{C}$ of a point $\mathbf{p}_{i}$ by bilinear interpolating on $\mathbf{B}$ using its projection to the BEV. Each point feature is then decoded into the point’s scene flow $\mathbf{d}_{i}\in\mathbb{R}^{3}$ representing how much this point needs to offset to rectify the shadow effect. After being rectified, the new point cloud is used to scatter the set of point features $\left\{\mathbf{f}_{i}\right\}$ to the BEV to obtain a new BEV image $\mathbf{B}^{\prime}$. While undergoing minimized impact of the shadow effect thanks to the rectification, $\mathbf{B}^{\prime}$ is sparse as only occupied pillars have features, which has a negative effect on detection accuracy [24]. On the other hand, $\mathbf{B}$ is semi-dense thanks to the occupancy leakage caused by the convolutions in the detection models’ backbone but possesses feature misalignment due to the shadow effect. To utilize the best of the two representations, they are fused before being fed to the Region Proposal Networks (RPN) for 3D bounding boxes. The complete pipeline is summarized in Fig.6.
Figure 6: Aligner
@@
To improve the accuracy of [1] in estimating scene flow, thus ultimately improving object detection, we introduce two extensions (i) incorporating HD Map and (ii) distilling an Oracle model. Previous works in integrating HD Maps into 3D object detection models ([26, 27]) opt for a mid-fusion approach that concatenates rasterized maps with backbone-made BEV representations. However, this approach is incompatible with copy-paste data augmentation [25], which randomly samples ground truth objects from a database and pastes them to each point cloud used for training, as pasted objects do not necessarily adhere to the semantics and geometry of the map (e.g., cars appear inside a building). Moreover, they only utilize the map’s binary channels such as drivable areas, sidewalks, or car parks while omitting the lane direction which is potentially helpful for estimating scene flow and objects’ heading direction.
@@
Figure 7: Extracting map features for points.
Figure 8: Alinger++ ’s Teacher-student framework
To improve the quality of the fused BEV representation, $\mathbf{B}^{f}$ in Fig.6, in terms of minimizing the feature misalignment caused by the shadow effect and reducing the sparsity, we use the teacher-student framework shown in Fig.8. In detail, we use ground truth objects’ trajectories to rectify the concatenation of point cloud sequences and use the shadow-effect-free results, illustrated in Fig.9, to train a teacher model. After the teacher converged, its BEV representation $\mathbf{B}^{t}$ is used to guide the student’s fused BEV representation $\mathbf{B}^{f}$ by optimizing the students’ weight so that the difference measured by the $\mathit{L}_{2}$ loss between these two representations is minimized.
Figure 9: Comparison between concatenation of a point cloud sequence (a) and its rectification using objects’ ground truth trajectories (b). Points are color-coded according to their time-lag with respect to the present. The color bar is shown in (c).
@@ III-B V2X Collaborative Perception Framework
The V2X setting that we target in this paper comprises multiple CAVs and IRSUs which are collectively referred to as agents. An agent $\mathcal{A}_{i}$ is equipped with a LiDAR to localize in a common global frame $\mathcal{G}$ and detect objects in its surrounding environment. The detection is based on the processing of point cloud sequences using a detection model made of the integration of the Aligner++ presented in the Sec.III-A2 into an off-the-shelf single-frame object detector such as PointPillar [22]. At a time step $t_{i}$, agent $\mathcal{A}_{i}$ uses a $K$-point-cloud sequence $\mathcal{S}_{i}=\left\{\mathcal{P}^{t_{i}-K+1}_{i},\dots,\mathcal{P}^{t_{i}}_{i}\right\}$ as an input to its detection model. An object $b_{i,j}$ detected by agent $\mathcal{A}_{i}$ is parameterized by a nine-vector $[x,y,z,w,l,h,\theta,s,c]$. The first seven numbers localize the object by its center location $[x,y,z]$, size $[w,l,h]$, heading direction $\theta$. The last two numbers $s$ and $c$ respectively denote confidence score $s$ and the predicted class $c$.
@@ III-B1 MoDAR for Object Detection on Point Cloud Sequences
MoDAR is created to enable detecting objects in extremely long point cloud sequences (hundreds of frames). In the MoDAR framework, a sequence is divided into several short sequences where objects are detected by a single-frame detector, tracked by a simple multi-object tracktor (e.g., [28]). Then, the data-driven motion forecasting model MultiPath++[29] predicts objects’ future poses based on objects’ trajectories established by the tracktor. Using prediction about objects’ future poses, detected objects in each short subsequence are propagated to the desired time step (e.g., the present). Next, each propagated object, which is represented as an up-right 3D bounding box (parameterized by the location of its center, size, and heading) with a confidence score and a class, is interpreted into a 3D point that takes the box’s center as its coordinate and the box’s size, heading, confidence score and class as features. These points interpreted from 3D bounding boxes are referred to as MoDAR points. They enable packing an entire subsequence into a small number of points, thus enabling an efficient fusion of extremely long point cloud sequences.
@@ III-B2 V2X Collaboration using MoDAR Points
The challenge in our V2X setting is that different agents detect objects at different rates, thus forcing the ego vehicle to utilize MoDAR points made by other agents at passed time steps. This timestamp mismatch results in a spatial misalignment between exchanged MoDAR points and ground truth dynamic objects, which can diminish the benefit of collaboration or even decrease the ego vehicle’s accuracy. This challenge is encountered in the context of single-vehicle detection on point cloud sequences as well because dynamic objects change their poses from one subsequence to another. [9] resolves this by predicting objects’ pose using a multi-object tracktor and the motion forecasting model MultiPath++.
While this is feasible in the V2X context, the implementation of add-on modules for future pose prediction does not align with our design target of minimal architecture. Instead, we use the scene flow to propagate MoDAR points from a passed timestep to the timestep queried by the ego vehicle. Since MoDAR points are virtual, their scene flow is not estimated directly from a point cloud sequence but is aggregated from the scene flow of points residing in the box they represent. To be concrete, let $\mathbf{m}_{i,j}$ be a MoDAR point representing a 3D bounding box $\mathrm{B}_{i,j}$ detected by agent $\mathcal{A}_{i}$ at timestep $t_{i}$. $\mathbf{p}_{i,h}$ and $\mathbf{f}_{i,h}$ respectively denote a real 3D point in the concatenation of agent $\mathcal{A}_{i}$’s point cloud sequence $\mathcal{S}_{i}$ and its predicted scene flow. The scene flow $\mathbf{f}^{\mathbf{m}}_{i,j}$ of $\mathbf{m}_{i,j}$ is computed by
$\mathbf{f}^{\mathbf{m}}_{i,j}=\mathrm{mean}\left\{\mathbf{f}_{i,h}~{}|~{}\mathbf{p}_{i,h}\in\mathrm{B}_{i,j}\right\}$ (4)
@@ IV-A1 Datasets and Metrics
As point clouds’ resolution exerts a large impact on the performance of LiDAR-based models, we test our model on three resolutions: 16, 32, and 64. While 32-channel and 64-channel point clouds are readily available in NuScenes dataset [2] and KITTI dataset [17] respectively, to the best of our knowledge, there are no publicly available datasets containing 16-channel point clouds. As a result, we synthesize a 16-channel dataset from the NuScenes dataset using the downsampling approach of [30].
@@
’s 3D Object Detection partition contains 7481 and 7581 samples for training and testing. Each sample comprises sensory measurements collected by a 64-beam LiDAR operating at 10Hz and several cameras. A common practice when working with KITTI is to split the original training data into 3712 training samples and 3769 validation samples for experimental studies. A challenge we encounter when using KITTI is that its 3D Object Detection partition contains temporally disjointed samples, thus being not straightforward to obtain input (point cloud sequences) and ground truth (scene flow) for our model. We resolve this challenge by matching samples in the 3D Object Detection partition with KITTI’s raw sequences. Since there are a few raw sequences that do not have tracklet annotation, meaning objects are not tracked, we only retain data samples whose associated raw sequences have tracklets (i.e., trajectories of objects) in the training set and validation set.
@@
We use mean Average Precision (mAP) to measure the performance of our model on the 3D object detection task. A prediction is matched with the ground truth that is closest, measured by an affinity, to it. A match is considered valid if the affinity is below a predefined threshold. For each threshold, the average precision is obtained by integrating the recall-precision curve for recall and precision above 0.1. The mAP is the mean of the average precision of the threshold set. For NuScenes, the affinity is the Euclidean distance on the ground plane between centers of predictions and ground truth. This distance has four thresholds: 0.5, 1.0, 2.0, and 4.0. For KITTI, the affinity is the Intersection-over-Union (IoU) in the BEV plane. Unlike NuScenes, KITTI uses only one affinity threshold for each class which is 0.7 for cars and 0.5 for pedestrians.
@@ IV-A2 Experiments and Results
The implementation of the single-vehicle perception model here follows the implementation made in our previous work [1] which uses the concatenation of 0.5-second point cloud sequences by EMC as the input. Since NuScenes and KITTI respectively obtain point clouds at 20 and 10 Hz, a 0.5-second sequence contains 10 and 5 point clouds. We use the same architecture, which uses PointPillar [22] as its backbone and CenterHead [4] as its RPN, for every experiment. In experiments on NuScenes, the detection range is limited to $[-51.2,51.2]$ on the XY plane and $[-5.0,3.0]$ along the Z axis. For experiments in KITTI, point clouds are first cropped using the left-color camera’s field of view as the annotations are only provided for objects that lie within the left-color camera’s image. The detection range on KITTI dataset is $[0,69.12]\times[-39.68,39.68]\times[-3,1]$ along X, Y, and Z axis. Our implementation is based on the OpenPCDet [32]. The details on the model’s hyperparameters can be found in our code release.
The evaluation of the Aligner++ on scene flow estimation task is done using the NuScenes dataset. The result shown in Tab.I indicates a significant improvement compared to our previous work [1] that reaches a state-of-art on accuracy-related metrics namely AccS, AccR, and ROutliers while maintaining low inference time. However, we have relatively high EPE this is because the evaluation is done for every point in the point cloud. This means the evaluation is carried out for both ground truth foreground points which have associated ground truth scene flow (can be nonzero or zero depending on dynamic) and ground truth background points which we assign the null vector as ground truth scene flow. Since the classification module of the Object Head inevitably makes false positive/ false negative foreground predictions, a number of background/ foreground points are predicted to have nonzero/ zero scene flow. Even though the portion of false predictions is small, as indicated by accuracy (AccS, AccR) and outlier metrics, the magnitude of their error is sufficiently large, due to the fact either the prediction or ground truth is zero, thus resulting in a large EPE.
@@
Aligner [1] 0.506 16.8 30.2 33.8 0.09
Aligner++ 0.616 46.4 66.6 6.8 0.10
The better scene flow estimation made by the results in a higher detection accuracy on the NuScenes dataset which can be seen in Tab.II. The integration of Aligner++, made of HD Map and distillation using the teacher-student framework, improves the mAP average over 10 classes of objects of a plain PointPillar by 6 points which is almost double the gain brought by Aligner (3.4 points). Interestingly, the distillation is responsible for most (5.6 out of 6 points) of the success of the Aligner++.
@@
The robustness of our Aligner++ is demonstrated by experiments on the KITTI and the synthetic NuScenes-16 dataset. As can be seen Tab.III and Tab.IV, the integration of Aligner++ consistently improves the accuracy of detecting objects in point cloud sequences. A common point between these two tables is that performance gain amounts to the Aligner++ is larger for vehicle-like classes (e.g., car, truck, or trailer). This is because the scene flow ground truth is generated using the rectification transformation (2) which is established based on the rigid motion assumption. This assumption holds for vehicle-like classes while being an oversimplification for pedestrians. As a result, the estimation of scene flow for vehicle-like classes is more accurate, thus higher detection accuracy. Another important result can be drawn from Tab.III and Tab.IV is that the effectiveness, in terms of detection accuracy, of using point cloud sequences over single point clouds persists on various LiDAR resolutions.

### v3

@@ III-A Learning Scene Flow
As mentioned in Sec.I, the accuracy of single-vehicle detection can be significantly improved by using concatenated point cloud sequences as inputs. However, such concatenation is done by the Ego Motion Compensation (EMC) method which only accounts for the motion of the ego vehicle resulting in the shadow effect, as can be seen in Fig.5, in the concatenated point cloud. This effect, which is a misalignment in the 3D space between objects’ points and their locations, leads to a misalignment in feature spaces (e.g., BEV representations), thus degrading detection accuracy for dynamic objects. In this section, we recall the approach we use in [17] to handle this effect and present the extensions we make to it.
Figure 5: A dynamic car’s appearance on the concatenation of a 10-point-cloud sequence spanning 0.5 seconds. Points are color-coded according to their time lag with respect to the present.
III-A1 The Cause of The Shadow Effect
@@
Let $\mathcal{P}^{t}=\left\{\mathbf{p}_{j}^{t}=[x,y,z,r,\Delta]~{}|~{}j=1,\dots,N\right\}$ denote a point cloud collected by the ego vehicle at time step $t$. Here, each point $\mathbf{p}_{j}$ has two features namely intensity $r$ and time-lag $\Delta$ with respect to a predefined time step. The point cloud is expressed in the ego vehicle frame $\mathcal{E}(t)$ measured with respect to a global frame $\mathcal{G}$. To concatenate a point cloud sequence $\mathcal{S}=\left\{\mathcal{P}^{t-K},\mathcal{P}^{t-K+1},\dots,\mathcal{P}^{t}\right\}$ of length $K+1$, each point of point cloud $\mathbf{P}^{t-\Delta t}$ is transformed to the global frame using the ego vehicle pose at the same time step as in (1), hence the method’s name Ego Motion Compensation.
@@
Here, ${}^{d}\mathbf{T}_{s}\in\mathbf{SE(3)}$ represents the rigid transformation that maps points in frame $s$ to frame $d$. EMC undoes the motion of the ego vehicle without accounting for objects’ motion. As a result, the appearance of a dynamic object in the concatenated point cloud comprises several instances, each corresponding to the object’s poses in the global frame at a particular time step, as in Fig.5.
@@ III-A2 Learning to Rectify the Shadow Effect
In [17], we propose a drop-in module, named Aligner, for 3D object detection models that creates a BEV representation, $\mathbf{B}\in\mathbb{R}^{C\times H\times W}$, of the scene as an intermediate output. Our module first computes the feature of $\mathbf{f}_{i}\in\mathbb{R}^{C}$ of a point $\mathbf{p}_{i}$ by the bilinear interpolation of $\mathbf{B}$ using its projection to the BEV. Each point feature is then decoded into the point’s scene flow $\mathbf{d}_{i}\in\mathbb{R}^{3}$ representing how much this point needs to offset to rectify the shadow effect. After being rectified, the new point cloud is used to scatter the set of point features $\left\{\mathbf{f}_{i}\right\}$ to the BEV to obtain a new BEV image $\mathbf{B}^{\prime}$. While undergoing minimized impact of the shadow effect thanks to the rectification, $\mathbf{B}^{\prime}$ is sparse as only occupied pillars have features, which has a negative effect on detection accuracy [25]. On the other hand, $\mathbf{B}$ is semi-dense thanks to the occupancy leakage caused by the convolutions in the detection models’ backbone but possesses feature misalignment due to the shadow effect. To utilize the best of the two representations, they are fused before being fed to the Region Proposal Networks (RPN) for 3D bounding boxes. The complete pipeline is summarized in Fig.6.
Figure 6: Aligner’s overall pipeline. Point-wise scene flow is predicted by Point Head based on point-wise features which are interpolated from the BEV image of the input point cloud. The input point cloud is then rectified by adding the corresponding scene flow to the coordinate of each LiDAR point. The scattering of point-wise features to the BEV using their rectified coordinate results in the rectified BEV image which is shadow-effect-free but sparse. The original BEV image which suffers from the shadow effect and the rectified BEV image are fused to minimize the shadow effect and sparsity. The resulting fused BEV image is finally fed to the RPN to obtain object detection.
@@
III-A3 Aligner++
To improve the accuracy of [17] in estimating scene flow, thus ultimately improving object detection, we introduce two extensions (i) incorporating HD Map and (ii) distilling a model trained on the concatenation of point cloud sequences by the ground truth trajectories.
Previous works in integrating HD Maps into 3D object detection models [26, 27] opt for a mid-fusion approach that concatenates rasterized maps with backbone-made BEV representations. However, this approach is incompatible with copy-paste data augmentation [12], which randomly samples ground truth objects from a database and pastes them to each point cloud used for training, as pasted objects do not necessarily adhere to the semantics and geometry of the map (e.g., cars appear inside a building). Moreover, they only utilize the map’s binary channels such as drivable areas, sidewalks, or car parks while omitting the lane direction which is potentially helpful for estimating scene flow and objects’ heading direction.
@@ III-A3 Aligner++
Figure 7: Extraction of map features. LiDAR points are projected to the BEV. Their map feature is obtained by nearest-neighbor interpolation on the rasterized HD Map.
Figure 8: Alinger++ ’s teacher-student framework. A model, pre-trained on point clouds concatenated using object ground truth trajectories, plays the role of the teacher. During training, the student model, which operates on point clouds concatenated using ego vehicle localization, strives to emulate the BEV images made by the teacher by minimizing a distillation loss.
To improve the quality of the fused BEV representation, $\mathbf{B}^{f}$ in Fig.6, in terms of minimizing the feature misalignment caused by the shadow effect and reducing the sparsity, we use the teacher-student framework shown in Fig.8. In detail, we use ground truth objects’ trajectories to rectify the concatenation of point cloud sequences and use the shadow-effect-free results, illustrated in Fig.9, to train a teacher model. After the teacher converges, its BEV representation $\mathbf{B}^{t}$ is used to guide the student’s fused BEV representation $\mathbf{B}^{f}$ by optimizing the students’ weight so that the difference measured by the $\mathit{L}_{2}$ loss between these two representations is minimized.
Figure 9: Comparison between concatenation of a point cloud sequence (a) and its rectification using objects’ ground truth trajectories (b). Points are color-coded according to their time lag with respect to the present. The color bar is shown in (c).
@@ III-B V2X Collaborative Perception Framework
The V2X setting that we target in this paper comprises multiple CAVs and IRSUs which are collectively referred to as agents. An agent $\mathcal{A}_{i}$ is equipped with a LiDAR to localize in a common global frame $\mathcal{G}$ and detect objects in its surrounding environment. The detection is based on the processing of point cloud sequences using a detection model made of the integration of the Aligner++ presented in the Sec.III-A2 into an off-the-shelf single-frame object detector such as PointPillar [23]. At a time step $t_{i}$, agent $\mathcal{A}_{i}$ uses a $K$-point-cloud sequence $\mathcal{S}_{i}=\left\{\mathcal{P}^{t_{i}-K+1}_{i},\dots,\mathcal{P}^{t_{i}}_{i}\right\}$ as an input to its detection model. An object $\mathbf{b}_{i,j}$ detected by agent $\mathcal{A}_{i}$ is parameterized by a nine-vector $[x,y,z,w,l,h,\theta,s,c]$. The first seven numbers localize the object by its center location $[x,y,z]$, size $[w,l,h]$, heading direction $\theta$. The last two numbers, $s$ and $c$, respectively denote confidence score $s$ and the predicted class $c$.
@@ III-B1 MoDAR for Object Detection on Point Cloud Sequences
MoDAR is created to enable detecting objects in extremely long point cloud sequences (hundreds of frames). In the MoDAR framework, a sequence is divided into several short sequences where objects are detected by a single-frame detector, tracked by a simple multi-object tracker (e.g., [16]). Then, the data-driven motion forecasting model MultiPath++[15] predicts objects’ future poses based on objects’ trajectories established by the tracker. Using prediction about objects’ future poses, detected objects in each short subsequence are propagated to the desired time step (e.g., the present). Next, each propagated object, which is represented as an up-right 3D bounding box (parameterized by the location of its center, size, and heading) with a confidence score and a class, is converted into a 3D point that takes the box’s center as its coordinate and the box’s size, heading, confidence score and class as features. These points are referred to as MoDAR points. They enable packing an entire subsequence into a small number of points, thus enabling an efficient fusion of extremely long point cloud sequences.
@@ III-B2 V2X Collaboration using MoDAR Points
The challenge in our V2X setting is that different agents detect objects at different rates, thus forcing the ego vehicle to utilize MoDAR points made by other agents at passed time steps. This timestamp mismatch results in a spatial misalignment between exchanged MoDAR points and ground truth dynamic objects, which can diminish the benefit of collaboration or even decrease the ego vehicle’s accuracy. This challenge is encountered in the context of single-vehicle detection on point cloud sequences as well because dynamic objects change their poses from one subsequence to another. [24] resolves this by predicting objects’ pose using a multi-object tracker and the motion forecasting model MultiPath++.
While this is feasible in the V2X context, the implementation of those modules for future pose prediction does not align with our design target of minimal architecture. Instead, we use the scene flow to propagate MoDAR points from a prior timestep to the timestep queried by the ego vehicle. Since MoDAR points are virtual, their scene flow is not estimated directly from a point cloud sequence but is aggregated from the scene flow of points residing in the box they represent. Specifically, let $\mathbf{m}_{i,j}$ be a MoDAR point representing a 3D bounding box $\mathbf{b}_{i,j}$ detected by agent $\mathcal{A}_{i}$ at timestep $t_{i}$. $\mathbf{p}_{i,h}$ and $\mathbf{f}_{i,h}$ respectively denote a real 3D point in the concatenation of agent $\mathcal{A}_{i}$’s point cloud sequence $\mathcal{S}_{i}$ and its predicted scene flow. The scene flow $\mathbf{f}^{\mathbf{m}}_{i,j}$ of $\mathbf{m}_{i,j}$ is computed by
$\mathbf{f}^{\mathbf{m}}_{i,j}=\mathrm{mean}\left\{\mathbf{f}_{i,h}~{}|~{}\mathbf{p}_{i,h}\in\mathbf{b}_{i,j}\right\}$ (4)
@@ IV Experiments and Results
This section first shows the results of the improvements to single vehicle perception with scene flow using real-world datasets, then demonstrates the collaborative approach using a simulated V2X dataset.
@@ IV-A1 Datasets and Metrics
As point clouds’ resolution has a large impact on the performance of LiDAR-based models, we test our model on three resolutions: 16, 32, and 64. While 32-channel and 64-channel point clouds are readily available in NuScenes dataset [6] and KITTI dataset [18] respectively, to the best of our knowledge, there are no publicly available datasets containing 16-channel point clouds. As a result, we synthesize a 16-channel dataset from the NuScenes dataset using the downsampling approach of [28].
@@
has a 3D Object Detection partition containing 7481 and 7581 samples for training and testing. Each sample comprises sensor measurements collected by a 64-beam LiDAR operating at 10Hz and several cameras. A common practice when working with KITTI is to split the original training data into 3712 training samples and 3769 validation samples for experimental studies. A challenge we encounter when using KITTI is that its 3D Object Detection partition contains temporally disjointed samples, thus being not straightforward to obtain input (point cloud sequences) and ground truth (scene flow) for our model. We resolve this challenge by matching samples in the 3D Object Detection partition with KITTI’s raw sequences. Since there are a few raw sequences that do not have tracklet annotation, meaning objects are not tracked, we only retain data samples whose associated raw sequences have tracklets (i.e., trajectories of objects) in the training set and validation set.
@@
We use mean Average Precision (mAP) to measure the performance of our model on the 3D object detection task. A prediction is matched with the closest ground truth, measured by an affinity. A match is considered valid if the affinity is below a predefined threshold. For each threshold, the average precision is obtained by integrating the recall-precision curve for recall and precision above 0.1. The mAP is the mean of the average precision of the threshold set. For NuScenes, the affinity is the Euclidean distance on the ground plane between centers of predictions and ground truth. This distance has four thresholds: 0.5, 1.0, 2.0, and 4.0. For KITTI, the affinity is the Intersection-over-Union (IoU) in the BEV plane. Unlike NuScenes, KITTI uses only one affinity threshold for each class, which is 0.7 for cars and 0.5 for pedestrians.
@@ IV-A2 Experiments and Results
The implementation of the single-vehicle perception model here follows the implementation made in our previous work [17] which uses the concatenation of 0.5-second point cloud sequences by EMC as the input. Since NuScenes and KITTI respectively obtain point clouds at 20 and 10 Hz, a 0.5-second sequence contains 10 and 5 point clouds. We use the same architecture, which uses PointPillar [23] as its backbone and CenterHead [8] as its RPN, for every experiment. In experiments on NuScenes, the detection range is limited to $[-51.2,51.2]$ on the XY plane and $[-5.0,3.0]$ along the Z axis. For experiments in KITTI, point clouds are first cropped using the left-color camera’s field of view as the annotations are only provided for objects that lie within the left-color camera’s image. The detection range on KITTI dataset is $[0,69.12]\times[-39.68,39.68]\times[-3,1]$ along the X, Y, and Z axis. Our implementation is based on the OpenPCDet [30]. The details on the model’s hyperparameters can be found in our code release.
Scene Flow Evaluation
The evaluation of the Aligner++ on scene flow estimation is done using the NuScenes dataset. The result shown in Tab.I indicates a significant improvement compared to our previous work [17] that reaches state-of-the-art on accuracy-related metrics namely AccS, AccR, and ROutliers while maintaining low inference time. The runtime of our methods, Aligner and Aligner++, is measured on an NVIDIA A6000 GPU, while the runtime of five baselines is measured on an NVIDIA RTX 3090 GPU (as reported by [31]) having the same compute capability. The Aligner++ has relatively high EPE; this is because the evaluation is done for every point in the point cloud. This means the evaluation is carried out for both ground truth foreground points which have associated ground truth scene flow (can be nonzero or zero depending on dynamic) and ground truth background points which we assign the null vector as ground truth scene flow. Since the classification module of the Object Head inevitably makes false positive/ false negative foreground predictions, a number of background/ foreground points are predicted to have nonzero/ zero scene flow. Even though the portion of false predictions is small, as indicated by accuracy (AccS, AccR) and outlier metrics, the magnitude of their error is sufficiently large, due to the fact either the prediction or ground truth is zero, thus resulting in a large EPE.
@@
PointPillar + 0.506 16.8 30.2 33.8 0.09
Aligner [17]
PointPillar + 0.616 46.4 66.6 6.8 0.10
Aligner++
Object Detection Evaluation
The better scene flow estimation made by the results in a higher detection accuracy on the NuScenes dataset which can be seen in Tab.II. The integration of Aligner++, made using the HD Map and through distillation using the teacher-student framework, improves the mAP average over 10 classes of objects of a plain PointPillar by 6 points which is almost double the gain brought by Aligner (3.4 points). Interestingly, the distillation is responsible for most (5.6 out of 6 points) of the success of the Aligner++.
@@
The robustness of our Aligner++ is demonstrated by experiments on the KITTI and the synthetic NuScenes-16 dataset. As can be seen Tab.III and Tab.IV, the integration of Aligner++ consistently improves the accuracy of detecting objects in point cloud sequences. A common point between these two tables is that performance gain from Aligner++ is larger for vehicle-like classes (e.g., car, truck, or trailer). This is because the scene flow ground truth is generated using the rectification transformation (2) which is established based on the rigid motion assumption. This assumption holds for vehicle-like classes while being an oversimplification for pedestrians. As a result, the estimation of scene flow for vehicle-like classes is more accurate, thus resulting in higher detection accuracy. Another important result that can be drawn from Tab.III and Tab.IV is the effectiveness, in terms of detection accuracy, of using point cloud sequences over single point clouds persists across various LiDAR resolutions.